OpenStudy (lena772):

If 3.85 moles of aluminum (Al) metal, with a density of 2.70 g/mL, are needed for a chemical reaction, how many milliliters of aluminum are needed? 1.43 mL Al 18.9 mL Al 38.5 mL Al 281 mL Al

OpenStudy (lena772):

@Luigi0210 @iPwnBunnies @Preetha

OpenStudy (ipwnbunnies):

Convert the moles of Aluminum to grams, using the molar mass of aluminum found on the Periodic Table. Since you have the density of aluminum, you can use dimensional analysis to find volume in mL.

OpenStudy (lena772):

3.85 mol Al * 26.98g Al / 1 mol Al =103.873g Al d=m/v vd=m v=m/d v=103.873g Al/2.7g/mL =38.47 =38.5 mL

OpenStudy (ipwnbunnies):

Mhm. c: